# LLM-Driven Usefulness Labeling for IR Evaluation

### v1

@@ Abstract.
In the information retrieval (IR) domain, evaluation plays a crucial role in optimizing search experiences and supporting diverse user intents. In the recent LLM era, research has been conducted to automate document relevance labels, as these labels have traditionally been assigned by crowd-sourced workers - a process that is both time and consuming and costly. This study focuses on LLM-generated usefulness labels, a crucial evaluation metric that considers the user’s search intents and task objectives, an aspect where relevance falls short. Our experiment utilizes task-level, query-level, and document-level features along with user search behavior signals, which are essential in defining the usefulness of a document. Our research finds that (i) pre-trained LLMs can generate moderate usefulness labels by understanding the comprehensive search task session, (ii) pre-trained LLMs perform better judgement in short search sessions when provided with search session contexts. Additionally, we investigated whether LLMs can capture the unique divergence between relevance and usefulness, along with conducting an ablation study to identify the most critical metrics for accurate usefulness label generation. In conclusion, this work explores LLM-generated usefulness labels by evaluating critical metrics and optimizing for practicality in real-world settings.
@@ 1. Introduction
Assessing relevance is not enough for many IR scenarios, as in many cases it cannot fully capture the extent to which a document is actually useful for addressing the motivating tasks. Although many scholars have advocated for this over decades (Cole et al., 2009; Liu, 2022), mainstream IR has remained focused on the seemingly objective assessment of relevance, following the traditional Cranfield evaluation paradigm (Cormack et al., 1998). Relevance is often used as a proxy for other constructs due to its strong correlation with user-side evaluation labels like usefulness and satisfaction (Mao et al., 2016a). However, usefulness is tied to both task context and users’ context-dependent preferences, making it a richer but more challenging construct to capture in real time. For Web search evaluations, relevance-based metrics alone cannot capture the user’s search goal and satisfaction, especially during complex task-driven sessions (Belkin et al., 2008; Liu, 2021). Therefore, incorporating usefulness as an evaluation metric provides a more holistic assessment, capturing not only document relevance, but also user search intent and satisfaction throughout sessions (Mitsui et al., 2017; Liu et al., 2020).
@@
To mitigate time and costly111As a concrete example, for the 50 topics in the TREC-8 Ad Hoc track (Voorhees and Harman, 2000), 129 participating systems led to more than 86,000 pooled documents to judge, requiring more than 700 assessor hours at a cost of about USD 15,000 (Thomas et al., 2024). constraints of traditional evaluation challenges (Thomas et al., 2024), IR researchers have been exploring the possibility and techniques of automated search evaluation with LLMs (Faggioli et al., 2023; Chen et al., 2024; Alaofi et al., 2024). LLMs for automated evaluation (LLM4eval) have recently attracted growing attention in IR community because of their flexibility and increasing task-solving capabilities (Rahmani et al., 2024a). In LLM-automated evaluations, researchers have primarily focused on relevance labels, while user-centric usefulness metrics remain largely unexplored. As usefulness is a measure that directly reflects a document’s contribution to achieving user goals, in this paper we particularly focus on usefulness as the key metric for evaluating Web search along with relevance. To achieve this, we generate usefulness labels using LLMs and employ search session, task-level, query-level and document-level user metrics, aiming to better capture how users perceive the quality of search systems. We also seek to explore whether LLMs can capture the diverging relationship between relevance and usefulness. In some cases, a retrieved document may be relevant to the query but not necessarily useful to the user, or it may be irrelevant to the query yet highly valuable for completing the task. Being able to identify and address such situations can open up new possibilities for enhancing IR applications, especially in complex, open-ended tasks.
@@
Existing ML-based approaches for predicting usefulness often overlook the broader task and user contexts. This is the place where LLMs can step in with their are pre-trained existing knowledge to better understand user, task, and query context through appropriate prompting. The purpose of generating the usefulness labels is to gauge how well and accurately these LLMs can perform in automated usefulness judgment when given all relevant features (both online and offline) (Chen et al., 2017; Zhang et al., 2018; Liu et al., 2018). Specifically, our research seeks to answer the following three research questions (RQs):
@@ 2. Related Works
In LLM automated search engine evaluation, most of the research has solely focused on automating relevance labels (Abbasiantaeb et al., 2024; Faggioli et al., 2023; Rahmani et al., 2024b; Upadhyay et al., 2024a), where else user-centric usefulness metrics have not been explored yet. Faggioli et al. (Faggioli et al., 2023) were among the first to examine how LLMs can be used for relevance judgment, offering unique insights into human-LLM collaboration. Their work explored direct prompting strategies, highlighting both the potential and limitations of using LLMs to generate relevance labels. Another prominent work in LLM-automated relevance assessment by Thomas et al. (Thomas et al., 2024) asserted that LLMs can accurately predict searcher’s preferences and perform relevance labeling as accurately as human labelers. LLMs have also been used to fill relevance judgment gaps (Abbasiantaeb et al., 2024) and, to some extent, studied for utility judgment using retrieval-augmented generation (RAG) (Zhang et al., 2024; Salemi and Zamani, 2024). To address the limitations of Thomas et al. (Thomas et al., 2024), Upadhyay et al. (Upadhyay et al., 2024a) conducted large-scale relevance assessment on DL TREC tracks using UMBRELLA (Upadhyay et al., 2024b), an LLM-based relevance assessor for academic TREC-style evaluation. Most LLM-automated IR evaluations have focused solely on relevance labeling, but relevance alone is insufficient, as usefulness directly reflects user goal achievement (Belkin et al., 2008; Cole et al., 2009). In this work, we explore usefulness as a key metric for web search evaluation alongside relevance. We generate usefulness labels using LLMs, incorporating implicit user search behavior signals like CTR and dwell time (Chuklin et al., 2013; Chapelle et al., 2009; Jiang et al., 2015). Additionally, we examine whether LLMs can capture contextual relationships in search sessions by integrating relevance and satisfaction labels to assess search system quality while exploring the contrast between relevance and usefulness.
@@ 3. Methodology
Our main goal is to evaluate the extent to which LLMs can generate usefulness labels when given user search behavior signals along with task-level, query-level and document-level information. In real world scenarios when annotators are asked to generate usefulness labels, they perform the labeling for a specific query and task with entire session history taken into context. In our experiment, we follow the same methodology where LLMs are provided vital information about each user’s search session and it generates a usefulness label for each task and for all the clicked documents.
We conduct the experiments using our version (shown in Figure 1) of the widely used DNA prompt template from the work of Thomas et al. (Thomas et al., 2024) that uses LLMs to accurately predict searcher’s preferences by generating relevance labels. We conduct experiments on 6 different types of pre-trained LLMs from GPT (Achiam et al., 2023), Llama (Dubey et al., 2024) and DeepSeek (Liu et al., 2024) to evaluate the ability to generate usefulness labels. Additionally, we explore LLM’s ability to process the divergence cases of relevance-usefulness to capture the nuances of search session contexts. Finally, our research includes an ablation study of various search behavior metrics to identify the most effective ones for input into an LLM for optimizing accuracy while considering real-world constraints, such as token limitations, model sizes and associated costs.
@@ 3.1. Datasets
For this experiment, we used two datasets: THUIR KDD’19 (Mao et al., 2016b) and TianGong-QRef (Chen et al., 2021), with behavioral log statistics shown in Table 1. Both datasets include query satisfaction, session satisfaction, and usefulness labels, while only THUIR-KDD’19 provides relevance labels from external assessors. They capture user search interactions (CLICK, SCROLL, HOVER, MOVE) and session timestamps. From these, we extracted CTR, URL dwell time, query dwell time, and task dwell time per user. From the statistics, we can see that THUIR-KDD’19 contains longer search sessions with 9 tasks, whereas QRef features shorter, task-free sessions. The variety in these datasets provided a diverse foundation for our study, allowing us to explore real-world user-centric search behaviors across both structured task-driven sessions and more open-ended search interactions.
@@ 3.2. Baseline Data and Session Data
After curating all the additional metrics, we create our LLM dataset by treating each data point as an individual batch regardless of sessions and call them baseline data. All corresponding queries, along with the metrics for each clicked document within a user’s task, are grouped into a single batch, which we refer to as session data. The reason for aggregating all these information together is that in practical scenario, the overall session history matters when labeling usefulness for each clicked documents. For a particular task, users typically either find the required information from the first query and the corresponding clicked documents, successfully completing their search, or they need to reformulate their query to refine the search results and then locate the desired information. For this reason, to preserve the contextual information of a search session for a task, we formulate one batch as one task session for one user and provide it to the LLMs as a part of session data. Formulating each search session for each user in Table 1, we get around 447 sessions from THUIR-KDD’19 dataset and 2,024 sessions from QRef dataset.
@@ 4. Experimental Setup
We conducted our primary experiment on two datasets: (1) baseline data without session context and (2) session data with previous search context. Both included document-level metrics (e.g., URL, title, summary, SERP snippets, rank) and behavioral metrics (e.g., CTR, URL-query-task dwell time), along with relevance and satisfaction labels for a comprehensive view of user interactions. Using zero-shot prompting as shown in Figure 1, we asked LLMs to label clicked document usefulness: 3 = Very Useful, 2 = Fairly Useful, 1 = Somewhat Useful, 0 = Not Useful at All. We evaluated OpenAI’s GPT-3.5 Turbo, GPT-4o, and Meta’s latest open-source LLMs (Llama 3.3 70B-Instruct, Llama 3.2 3B-Instruct, Llama 3.1 8B-Instruct) (Dubey et al., 2024) for usefulness judgment. Given the task’s complexity, we included DeepSeek-R1-Distill-Qwen-14B for its strong reasoning. Experiments were conducted via AWS Bedrock for Llama models and OpenAI APIs for GPT models, using a temperature of 0 and top_p of 1. DeepSeek-R1-Distill-Qwen-14B was hosted locally via Ollama with the same temperature parameter.
@@ 5. Results
Our experiments address the three RQs (explained below) and evaluate the effectiveness of LLM-generated usefulness labels.
@@ 5.1. RQ1: Baseline and Session Analysis
We prompt the LLM models with our DNA prompt and generate usefulness labels for each of the clicked documents. Both datasets already contain the usefulness labels annotated by users that we use as our base ground truth. To address our RQ1, we evaluate how well the LLM-generated usefulness performs with respect to our ground truth. As the usefulness labels are ordinal, starting from 0 (low) to 3 (high), we conduct Spearman’s rank correlation () to measure the strength and direction of association between two ranked variables. The detailed analysis in Table 2 shows a moderate positive correlation between human and LLM-generated usefulness labels. Considering Baseline analysis, for THUIR-KDD’19 dataset, Llama 3.3 70B-Instruct has the highest correlation ( = 0.38), followed by GPT-3.5turbo ( = 0.37). The weakest correlation is observed in Llama 3.2 3B-Instruct ( = 0.19), which is likely due to its relatively small parameter size shown in Figure 2. For the QRef dataset, the overall trend is similar, with Llama 3.3 70B-Instruct performing moderately well ( = 0.33). Task-level and query-level findings for THUIR-KDD’19 align closely with overall trends, reinforcing consistency in our findings across all LLM models. In QRef’s baseline dataset, both GPT-4omini ( = 0.44) and Llama 3.3 70B-Instruct ( = 0.44) perform moderately well. For Session analysis, performance gradually improves from baseline, indicating that LLMs benefit from session-context search data. In QRef, particularly at the query level, a moderate improvement is observed in GPT models. This suggests that LLMs perform better in shorter (QRef) rather than longer (THUIR-KDD’19) search sessions. The best-performing models in both datasets are denoted with (+) and (*). Notably, in QRef session data at the query level ( = 0.48 and 0.47), GPT models perform the best. This further supports the observation that session-level context enhances LLM performance, particularly in query-level analysis. Despite DeepSeek’s proficient mathematical reasoning, it has struggled in this task. Instead of providing direct usefulness labels, it often generated reasoning-based responses, requiring additional extraction of the labels. This issue is particularly evident in cases where DeepSeek prioritizes generating justifications rather than concise label assignments. Llama 3.3 70B-Instruct remains the best-performing model across different levels of analysis. Its 70B parameter count likely contributes to its superior ability to process and interpret nuanced contextual information. GPT-4omini follows, performing well in both baseline and session contexts. Smaller models like Llama 3.2 3B-Instruct struggle to generate usefulness labels but improve in QRef’s session dataset at the query level ( = 0.46).
@@ 5.2. RQ2: Relevance vs Usefulness Divergence
We analyze the correlation between relevance and usefulness across four cases: HR-HU (High Relevance, High Usefulness), HR-LU (High Relevance, Low Usefulness), LR-HU (Low Relevance, High Usefulness), and LR-LU (Low Relevance, Low Usefulness). As only THUIR-KDD’19 dataset has relevance labels, we divide the dataset into the four cases of usefulness-relevance combinations and tested the performance of LLMs. From our initial observation, we report the Spearman rank correlation () of the baseline and session data for 5 selected LLMs in Table 3. To address our RQ2, we find that LLMs face challenges in capturing the contrasting relationship between relevance and usefulness in both baseline and session analysis. According to Spearman’s rank correlation (), all five LLM models are showing weak positive correlation and few of them have weak negative correlation. For the diverging cases HR-LU and LR-HU, the models face challenges in generating accurate usefulness labels.

### v2

@@ Abstract.
In the information retrieval (IR) domain, evaluation plays a crucial role in optimizing search experiences and supporting diverse user intents. In the recent LLM era, research has been conducted to automate document relevance labels. These labels have traditionally been assigned by crowd-sourced workers, a process that is both time consuming and costly. This study focuses on LLM-generated usefulness labels, a crucial evaluation metric that considers the user’s search intents and task objectives, an aspect where relevance falls short. Our experiment utilizes task-level, query-level, and document-level features along with user search behavior signals, which are essential in defining the usefulness of a document. Our research finds that (i) pre-trained LLMs can generate moderate usefulness labels by understanding the comprehensive search task session, and (ii) pre-trained LLMs perform better judgment in short search sessions when provided with search session contexts. Furthermore, we investigate whether LLMs can capture the unique divergence between relevance and usefulness, along with conducting an ablation study to identify the most critical metrics for accurate usefulness label generation. In conclusion, this work explores LLM-generated usefulness labels by evaluating critical metrics and optimizing for practicality in real-world settings.
@@ 1. Introduction
Assessing relevance is not enough for many IR scenarios, as in many cases it cannot fully capture the extent to which a document is actually useful for addressing the motivating tasks. Although many scholars have advocated this for decades (Cole et al., 2009; Liu, 2022), mainstream IR has remained focused on the seemingly objective assessment of relevance, following the traditional Cranfield evaluation paradigm (Cormack et al., 1998). Relevance is often used as a proxy for other constructs due to its strong correlation with user centric evaluation labels like usefulness and satisfaction (Mao et al., 2016a). However, usefulness is tied to both task context and users’ context dependent preferences, making it a richer but more challenging construct to capture in real time. For web search evaluations, relevance based metrics alone cannot capture the user’s search goal and satisfaction, especially during complex task driven sessions (Belkin et al., 2008; Liu, 2021). Therefore, incorporating usefulness as an evaluation metric provides a more holistic assessment, capturing not only document relevance, but also user search intent and satisfaction throughout sessions (Mitsui et al., 2017; Liu et al., 2020).
@@
To mitigate time and cost constraints of traditional evaluation challenges (Thomas et al., 2024), IR researchers have been exploring the possibility and techniques of automated search evaluation with LLMs (Faggioli et al., 2023; Chen et al., 2024; Alaofi et al., 2024). LLMs for automated evaluation (LLM4eval) have recently attracted growing attention in IR community because of their flexibility and increasing task-solving capabilities (Rahmani et al., 2024a). In LLM automated evaluations, researchers have primarily focused on relevance labels, while user centric usefulness metrics remain largely unexplored. As usefulness is a measure that directly reflects a document’s contribution to achieving user goals, in this paper we particularly focus on usefulness as the key metric for evaluating web search along with relevance. To achieve this, we generate usefulness labels using LLMs and employ search session, task-level, query-level and document-level user metrics, aiming to better capture how users perceive the quality of search systems. We also seek to explore whether LLMs can capture the diverging relationship between relevance and usefulness. In some cases, a retrieved document may be relevant to the query but not necessarily useful to the user, or it may be irrelevant to the query yet highly valuable for completing the task. Being able to identify and address such situations can open up new possibilities for enhancing IR applications, especially in complex, open ended tasks.
@@
Existing ML-based approaches for predicting usefulness often overlook the broader task and user contexts. This is the place where LLMs can step in with their pre-trained existing knowledge to better understand user, task, and query context through appropriate prompting. The purpose of generating usefulness labels is to gauge how well and accurately these LLMs can perform in automated usefulness judgment when given all relevant features (both implict and explicit) (Chen et al., 2017; Zhang et al., 2018; Liu et al., 2018; Fox et al., 2005). Specifically, our research seeks to answer the following three research questions (RQs):
@@ 2. Related Works
In LLM enabled search engine evaluation, most of the research has solely focused on automating relevance labels (Abbasiantaeb et al., 2024; Faggioli et al., 2023; Rahmani et al., 2024b; Upadhyay et al., 2024a), where else user centric usefulness metrics have not been explored yet. Faggioli et al. (Faggioli et al., 2023) were among the first to examine how LLMs can be used for relevance judgment, offering unique insights into human-LLM collaboration. Their work explored direct prompting strategies, highlighting both the potential and limitations of using LLMs to generate relevance labels. Another prominent work in LLM automated relevance assessment by Thomas et al. (Thomas et al., 2024) asserted that LLMs can accurately predict searcher’s preferences and perform relevance labeling as accurately as human labelers. Studies conclude that LLMs are more positive in assessing relevance and thus should be used to assess all documents, not just relevance gaps (Abbasiantaeb et al., 2024) and, to some extent LLMs are studied for utility judgment using retrieval-augmented generation (RAG) (Zhang et al., 2024; Salemi and Zamani, 2024). To address some limitations of Thomas et al. (Thomas et al., 2024), Upadhyay et al. (Upadhyay et al., 2024a) conducted large scale relevance assessment on DL TREC tracks using UMBRELLA (Upadhyay et al., 2024b), an LLM based relevance assessor for academic TREC-style evaluation. Most LLM automated IR evaluations have focused solely on relevance labeling, but relevance alone is insufficient, as usefulness directly reflects user goal achievement (Belkin et al., 2008; Cole et al., 2009). In this work, we explore usefulness as a key metric for web search evaluation alongside relevance. We generate usefulness labels using LLMs, incorporating implicit user search behavior signals like CTR and dwell time (Chuklin et al., 2013; Chapelle et al., 2009; Jiang et al., 2015). Additionally, we examine whether LLMs can capture contextual relationships in search sessions by integrating relevance and satisfaction labels to assess search system quality while exploring the contrast between relevance and usefulness.
@@ 3. Methodology
Our main goal is to evaluate the extent to which LLMs can generate usefulness labels when given user search behavior signals along with task-level, query-level and document-level information. In real-world scenarios, when evaluation is conducted by crowd-sourced workers they perform individual search sessions for each task and assign usefulness labels for specific query and task within the context of the entire search session. In our experiment, we follow the same methodology where LLMs are provided vital information about each user’s search session and it generates a usefulness label for each task and for all the clicked documents.
We conduct the experiments using our zero-shot prompt111https://github.com/moulydewan/LLMUsefulnessLabels that has been created taking inspiration from the structure of the widely used DNA prompt template by Thomas et al. (Thomas et al., 2024) that uses LLMs to generate relevance labels. We conduct experiments on 6 different types of pre-trained LLMs such as GPT-4omini, GPT-3.5turbo (Achiam et al., 2023), Llama 3.1 8B-Instruct, Llama 3.2 3B-Instruct, Llama 3.3 70B-Instruct (Dubey et al., 2024) and DeepSeek-R1-Distill-Qwen-14B (Liu et al., 2024) to evaluate the ability to generate usefulness labels. Additionally, we explore LLM’s ability to process the divergence cases of relevance-usefulness to capture the nuances of search session contexts. Finally, our research includes an ablation study of various search behavior metrics to identify the most effective ones for input into an LLM for optimizing accuracy while considering real-world constraints, such as token limitations, model sizes and associated costs.
@@ 3.1. Datasets
For this experiment, we used two datasets: THUIR KDD’19 (Mao et al., 2016b) and TianGong-QRef (Chen et al., 2021), with behavioral log statistics shown in Table 1. Both datasets include query satisfaction, session satisfaction, and usefulness labels while only THUIR-KDD’19 provides relevance labels from external assessors. They capture user search interactions (CLICK, SCROLL, HOVER, MOVE) and session timestamps. From these, we extracted CTR, URL dwell time, query dwell time, and task dwell time per user. From the statistics, we can see that THUIR-KDD’19 contains longer search sessions with 9 tasks, whereas QRef features shorter task-free sessions. For this reason, we perform task-level and query-level analysis for THUIR-KDD’19 and for QRef we perform session-level and query-level analysis (further shown in Table 2). The variety in these datasets provided a diverse foundation for our study, allowing us to explore real-world user-centric search behaviors across both structured task-driven sessions and more open ended task-free search interactions.
@@ 3.2. Baseline Data and Session Data
After curating all the additional metrics, we create our LLM dataset by treating each data point as an individual batch regardless of sessions and call them baseline data. Each data point represents a single user click on a web document with the associated metrics irrespective of user sessions. All corresponding queries, along with the metrics for each clicked document within a user’s task, are grouped into a single batch, which we refer to as session data. The reason for aggregating all these information together is that in practical scenario, the overall session history matters when labeling usefulness for each clicked documents. For a particular task, users typically either find the required information from the first query and the corresponding clicked documents, successfully completing their search, or they need to reformulate their query to refine the search results and then locate the desired information. If a user finds document X useful, a similar relevant document Y seen later may feel redundant and not useful. Yet without session context, the LLM might still rate both X and Y as equally useful. For this reason, to preserve the contextual information of a search session for a task, we formulate one batch as one task session for one user and provide it to the LLMs as a part of session data. Formulating each search session for each user in Table 1, we get around 447 sessions from THUIR-KDD’19 dataset and 2,024 sessions from QRef dataset.
@@ 4. Experimental Setup
We conduct our primary experiment on two datasets: (1) baseline data without session context and (2) session data with previous search context. Both include document-level metrics (e.g., URL, title, summary, SERP snippets, rank) and behavioral metrics (e.g., CTR, URL-query-task dwell time), along with relevance and satisfaction labels for a comprehensive view of user interactions. Using our zero-shot prompting, we asked LLMs to label clicked document usefulness: 3 = Very Useful, 2 = Fairly Useful, 1 = Somewhat Useful, 0 = Not Useful at All. We evaluate OpenAI’s GPT-3.5 Turbo, GPT-4omini, and Meta’s latest open-source LLMs (Llama 3.3 70B-Instruct, Llama 3.2 3B-Instruct, Llama 3.1 8B-Instruct) (Dubey et al., 2024) for usefulness judgment. For LLM coverage, we included the reasoning model DeepSeek-R1-Distill-Qwen-14B along with other LLMs. Experiments were conducted via AWS Bedrock for Llama models and OpenAI APIs for GPT models, using a temperature of 0 and top_p of 1. DeepSeek-R1-Distill-Qwen-14B was hosted locally via Ollama with the same temperature parameter.
@@ 5. Results
Our experiments address the three RQs for baseline and session data and evaluate LLM-generated usefulness at the overall level across all user interactions, while task and query scores are averaged separately to capture different granularity of ranking quality.
@@ 5.1. RQ1: Baseline and Session Analysis
We prompt the LLM models with our prompt and generate usefulness labels for each of the clicked documents. Both datasets already contain the usefulness labels annotated by humans that we use as our ground truth. To address our RQ1, we evaluate how well the LLM-generated usefulness performs with respect to our ground truth. As the usefulness labels are ordinal, starting from 0 (low) to 3 (high), we conduct Spearman’s rank correlation () to measure association between ranked variables. The detailed analysis in Table 2 shows a moderate positive correlation between human and LLM-generated usefulness labels. The best performing models in both datasets are denoted with + for baseline and * for session. Considering Baseline analysis, for THUIR-KDD’19 dataset for overall, Llama 3.3 70B-Instruct has the highest correlation ( = 0.38), followed by GPT-3.5turbo ( = 0.37). The weakest correlation is observed in Llama 3.2 3B-Instruct ( = 0.19), which is likely due to its relatively small parameter size (also shown in Figure 1). For the QRef dataset, the overall trend is similar with Llama 3.3 70B-Instruct performing moderately well ( = 0.33). Task-level and query-level findings for THUIR-KDD’19 align closely with overall trends, reinforcing consistency in our findings across all LLM models. In QRef’s baseline dataset for query level, both GPT-4omini ( = 0.44) and Llama 3.3 70B-Instruct ( = 0.44) perform the best. For Session analysis, performance gradually improves compared to baseline, indicating that LLMs benefit from session-context search data. In QRef, improvement is observed in session compared to baseline in overall, task-level and query-level session data. This suggests that LLMs perform better in shorter (QRef) rather than longer (THUIR-KDD’19) search sessions. Notably, in QRef session data at the query level ( = 0.48 and 0.47), GPT models perform the best among all. This further supports the observation that session-level context enhances LLM performance, particularly in query-level analysis. Despite DeepSeek’s proficient reasoning ability, it has struggled in this task. During usefulness label generation, even after direct prompting instead of providing direct usefulness labels it often generated reasoning-based responses, requiring additional filtering and extraction of the labels from the generated responses. From our experiments, Llama 3.3 70B-Instruct remains the best performing model across different levels of analysis. Its 70B parameter count likely contributes to its superior ability to process and interpret nuanced contextual information. GPT-4omini follows next, performing well in both baseline and session contexts. Smaller models like Llama 3.2 3B-Instruct struggle to generate usefulness labels but improve performance with session data particularly in QRef’s session data at both session ( = 0.37) and query level ( = 0.46).
@@ 5.2. RQ2: Relevance vs Usefulness Divergence
We analyze the correlation between relevance and usefulness across four cases: HR-HU (High Relevance, High Usefulness), HR-LU (High Relevance, Low Usefulness), LR-HU (Low Relevance, High Usefulness) and LR-LU (Low Relevance, Low Usefulness). As only THUIR-KDD’19 dataset has relevance labels we choose this dataset to conduct this part of experiment. We divide the dataset into four cases of usefulness-relevance combinations and tested the performance of LLMs. Both the relevance and usefulness labels are labeled from 0 to 3 where we consider 0 & 1 as Low and 3 & 4 as High. From our initial observation, we report the Spearman rank correlation () of the baseline and session data for 5 selected LLMs in Table 3. To address our RQ2, we find that LLMs face challenges in capturing the contrasting relationship between relevance and usefulness in both baseline and session analysis. According to Spearman’s rank correlation (), all five LLM models are showing weak positive correlation and few of them have weak negative correlation. For the diverging cases HR-LU and LR-HU, the models face challenges in generating accurate usefulness labels. This may be due to the nature of Spearman’s rank correlation, which relies on the overall rank order of data points. When the dataset is divided into four cases, the reduced range of rank variation and smaller sample sizes can weaken the correlation, making it harder to capture the relationship between human and LLM-generated usefulness labels within each subgroup. Furthermore, it shows that LLMs generalize measures such as high relevance will lead to high usefulness which is not the case always in real settings. The DNA structured prompt is solely not enough to capture the divergent cases, as such cases should be explicitly mentioned in the prompts for better clarification if expanded for future research.